# Stiefel-Whitney Classes for Finite Symplectic Groups

### v1

@@ 3.3. Product Formula for Total SWCs
Let $2^{a+1}$ be the highest power of $2$ which divides $|\mathcal{S}|$. Gorenstein in the proof of [Gor68, Theorem 8.3, Chapter 2] specifies elements $x,y\in\mathcal{S}$ which generate a $2$-Sylow subgroup of $\mathcal{S}$. (It is generalized quaternion of order $2^{a+1}$.) One checks that the subgroup $Q$ generated by $y$ and $x^{2^{a-2}}$ is the quaternion group of order $8$. As $X$ is isomorphic to $\mathcal{S}^{n}$, it correspondingly has a subgroup $Q_{X}$ isomorphic to $Q^{n}$, and containing $Z_{X}$. We thus have a chain of subgroups $Z_{X}\leq Q_{X}\leq X\leq G$. Since $Z_{X}$ detects SWCs of $G$, we infer that $Q_{X}$ also detects the SWCs of $G$.
@@ 3.4. Gow’s Formula
Theorem 3.8 ([Gow85], Theorem 1).
Let $\pi$ be an irreducible self-dual representation of $G$ with central character $\omega_{\pi}.$ Then, we have
@@
We simply call this equality Gow’s formula. For irreducible orthogonal $\pi$ of $G$, it means
$\chi_{\pi}(\mathbbm{1})=\chi_{\pi}(-\mathbbm{1}).$
This is same as $\chi_{\pi}(g_{0})=\chi_{\pi}(g_{n})$, which leads to
@@
$m_{k}(\pi)=\begin{cases}0&\text{ when }k\text{ is odd}\\ \cfrac{1}{2^{n-1}}\sum\limits_{i=0}^{\frac{n-1}{2}}\chi_{\sigma_{i}}(g_{k})\chi_{\pi}(g_{i})&\text{ when }k\text{ is even, }n\text{ is odd}\\[10.0pt] \cfrac{1}{2^{n-1}}\sum\limits_{i=0}^{\frac{n-2}{2}}\chi_{\sigma_{i}}(g_{k})\chi_{\pi}(g_{i})+\cfrac{1}{2^{n}}\big{(}\chi_{\sigma_{\frac{n}{2}}}(g_{k})\chi_{\pi}(g_{\frac{n}{2}})\big{)}&\text{ when }k,n\text{ both are even}.\end{cases}$
@@
$m_{k}(\varphi)=\begin{cases}0&\text{ when }k\text{ is even and }n\text{ is odd}\\ \cfrac{1}{2^{n-1}}\sum\limits_{i=0}^{\lceil\frac{n-2}{2}\rceil}\chi_{\sigma_{i}}(g_{k})\chi_{\varphi}(g_{i})&\text{ when }k\text{ is odd, }\\[10.0pt] \cfrac{1}{2^{n}}\chi_{\sigma_{\frac{n}{2}}}(g_{k})\chi_{\varphi}(g_{\frac{n}{2}})&\text{ when }k,n\text{ both are even}.\end{cases}$
@@ 4.1. The case $n=1$
When $\pi$ is irreducible orthogonal, $m_{1}(\pi)=0$ by Corollary 3.8.1. Therefore, $w(\pi)$ is trivial for such representations. (Compare [MS23, Theorem 1.2].)
@@ 4.2. The case $n=2$
(16) $\begin{split}w_{4}(\pi)&=\frac{m_{1}(\pi)+m_{2}(\pi)}{4}(\mathfrak{e}_{1}+\mathfrak{e}_{2})\\ &=\frac{1}{8}\left(\deg\pi-\chi_{\pi}(g_{1})\right)\mathcal{E}_{1}\text{, and}\\ \end{split}$
(17) $\begin{split}w_{8}(\pi)&=m_{1}(\pi)\mathcal{E}_{2}+\left(\binom{m_{1}(\pi)}{2}+\binom{m_{2}(\pi)}{2}\right)\mathcal{E}_{1}^{2}.\\ \end{split}$
@@
Next is a representation induced from $X$. Let $\psi$ be an odd linear character of $\mathbb{F}_{q^{2}}^{\times}$ with $\psi^{q}\neq\psi$. According to [FH13, Section 5.2], there is an irreducible cuspidal representation $\sigma_{\psi}$ of $\mathcal{S}$ corresponding to $\psi$, with degree $q-1$. Viewing $X$ as the product $\mathcal{S}\times\mathcal{S}$ as in Section 3.1, we form the external tensor product representation $\sigma_{\psi}\boxtimes 1$ of $X$. Finally, let $\pi_{2}=\operatorname{Ind}_{X}^{G}(\sigma_{\psi}\boxtimes 1)$. Here, $\deg(\pi_{2})=q^{2}(q-1)(q^{2}+1)$ and again as $\psi$ is odd, $\chi_{\pi_{2}}(-\mathbbm{1})=-\deg\pi_{2}.$ Also, $\chi_{\pi_{2}}(g_{1})=0$: observe that the only conjugates of $g_{1}$ in $X$ are ±$\pm g_{1}$ and there is a Weyl element sending $g_{1}$ to $-g_{1}$ upon conjugation. (And again use the Frobenius character formula.)
@@ 4.4. Weil Representations
defines a group isomorphism $\mathbb{F}_{q}\overset{\sim}{\to}\widehat{\mathbb{F}}_{q}$. Let $\alpha$ be a non-square, and $1$ be the identity in $\mathbb{F}_{q}^{\times}$. According to Gérardin [Gér77, Theorem 2.4(d)], the symplectic group $G=\operatorname{Sp}(2n,q)$ has two nonisomorphic Weil representations $\mathcal{W}_{\lambda_{1}}$, $\mathcal{W}_{\lambda_{\alpha}}$ associated to the characters $\lambda_{1}$, $\lambda_{\alpha}$. Set $\mathcal{W}=\mathcal{W}_{\lambda_{1}}$ and $\mathcal{W}^{\prime}=\mathcal{W}_{\lambda_{\alpha}}$. Both $\mathcal{W}$, $\mathcal{W}^{\prime}$ are complex representations of degree $q^{n}$, and decompose into two irreducible representations as
@@
From[Gér77, Corollary 4.8.1], the character values of $\mathcal{W}$ at $g_{i}$ are given by:
@@
where $[f_{k}]_{i}$ is the coefficient of $x^{i}$ in $f_{k}(x)=(1-x)^{k}(1+x)^{n-k}$. Note that $[f_{k}]_{n-i}=(-1)^{k}[f_{k}]_{i}$ for each $i=0,\ldots,n$, so that
@@
We have already shown this for $\mathcal{W}$. In [Gér77], we see that $\chi_{\mathcal{W}}(g_{i})=\chi_{\mathcal{W}^{\prime}}(g_{i})$ for each $i$. Thus, the formula also holds for $w(S(\mathcal{W}^{\prime}))$. ∎
@@ 5. Universal SWCs
This is implicit in [FP06, Theorem 6.1, page 283]. ∎
@@
By Lemma 3.2, it is enough to show that the restriction $\operatorname{H}^{i}(X_{m})^{S_{m}}\to\operatorname{H}^{i}(X_{n-1})^{S_{n-1}}$ is injective.
Write
$R_{n}=P_{n}\otimes_{\mathbb{F}_{2}}Q_{n}$
where $P_{n}=\mathbb{F}_{2}[\mathcal{E}_{1},\ldots,\mathcal{E}_{n}]$ is the subalgebra generated by $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$ and $Q_{n}=\mathbb{F}_{2}[\mathcal{F}_{1},\ldots,\mathcal{F}_{n}]$ is the subalgebra generated by $\mathcal{F}_{1},\ldots,\mathcal{F}_{n}$ in $\operatorname{H}^{*}(\mathcal{S}^{n})$.
From above, we need to show that the restriction of the map $P_{m}\otimes Q_{m}\to P_{n-1}\otimes Q_{n-1}$ to the degree $k$ part is injective for all $k<4n-1$.
@@
If $i^{*}_{n}(\alpha)=0$, then all $c_{\mathbf{r},\mathbf{s}}$ are zero since $\mathcal{E}^{\mathbf{r}}\mathcal{F}^{\mathbf{s}}$ are linearly independent. This implies $\alpha=0$, and hence we have the injectivity of $\iota^{*}_{n}$ for $k<4n-1$. ∎
@@ 6.1. The Polynomials $\operatorname{P}_{m,n}$
It is symmetric in the $x_{i}$ and the $y_{j}$ separately, so by the fundamental theorem of symmetric polynomials [Bou13, Chapter 4, Section 6, Theorem 1], there is a unique polynomial $\operatorname{P}_{m,n}\in\Lambda$ such that
@@
6.2. Characteristic Classes of External Tensor Products
@@
We will sketch the proof for SWCs in this Appendix; the case of CCs is similar. Iterating Proposition 6.1 gives:
Corollary 6.1.1.
@@
6.3. Products of Vector Bundles
@@
Proposition 6.2 ([MS16], Chapter 7, Problem 7-C).
@@
Consider the projection maps $p_{i}:B_{1}\times B_{2}\to B_{i}$. Let $p_{i}^{*}E_{i}$ be the pullback of $E_{i}$ by $p_{i}$ consisting of elements $((b_{1},b_{2}),e_{i})\in(B_{1}\times B_{2})\times E_{i}$ such that
$\Pi_{i}(e_{i})=p_{i}(b_{1},b_{2})=b_{i}.$
These are vector bundles over the same base space $B_{1}\times B_{2}$. Thus, we construct their internal tensor product, and put
$E_{1}\boxtimes E_{2}:=p_{1}^{*}E_{1}\otimes p_{2}^{*}E_{2},$
which is again a vector bundle over $B_{1}\times B_{2}.$ This with the naturality of SWCs and Proposition 6.2 gives:
Proposition 6.3.
@@
6.4. Vector Bundles From Representations
Let $G$ be a finite group, and $(\rho,U)$ be a real representation of $G$. Associated to $G$ is a classifying space $BG$ with a contractible right principal $G$-bundle $EG$. From $(\rho,U)$ one can form the associated real vector bundle $EG[U]$ over $BG$. To define $EG[U]$, first form the product $EG\times U$; it is $G$-space under the action $g\cdot(x,u)=(x\cdot g^{-1},\rho(g)u)$. Then $EG[U]$ is the quotient of $EG\times U$ by this action. Then put $w_{i}^{\mathbb{R}}(\rho)=w_{i}(EG[U])$. (See for instance [Ben98] or [GKT89].) The singular cohomology $\operatorname{H}^{*}(BG,A)$ is naturally isomorphic to the group cohomology $\operatorname{H}^{*}(G,A)$ for any abelian group $A$.
Let $(\pi,V)$ be a complex orthogonal representation of $G$. There is a representation $(\rho,U)$, with $U$ a real vector space, so that $\rho\otimes_{\mathbb{R}}\mathbb{C}\cong\pi$ (see [BD95, Chapter II, Section 6] for instance). Such a representation is called a real form of $\pi$; it is unique up to isomorphism. We now can define $w(\pi):=w^{\mathbb{R}}(\rho)$.
@@ 6.4. Vector Bundles From Representations
$\begin{split}w(\pi_{1}\boxtimes\pi_{2})&=w^{\mathbb{R}}(\rho_{1}\boxtimes\rho_{2})\\ &=w(EG[U_{1}\otimes U_{2}])\\ &=w(\mathcal{U}_{1}\boxtimes\mathcal{U}_{2})\in\operatorname{H}^{*}(BG_{1}\times BG_{2},\mathbb{F}_{2}),\\ \end{split}$
where $\mathcal{U}_{i}=EG_{i}[U_{i}]$. We may identify $BG_{1}\times BG_{2}$ with $BG$. Now, Proposition 6.1 for SWCs follows from Proposition 6.3 with $E_{i}=\mathcal{U}_{i}$.

### v2

@@ 3.3. Product Formula for Total SWCs
Let $2^{a+1}$ be the highest power of $2$ which divides $|\mathcal{S}|$. Gorenstein in the proof of [goren, Theorem 8.3, Chapter 2] specifies elements $x,y\in\mathcal{S}$ which generate a $2$-Sylow subgroup of $\mathcal{S}$. (It is generalized quaternion of order $2^{a+1}$.) One checks that the subgroup $Q$ generated by $y$ and $x^{2^{a-2}}$ is the quaternion group of order $8$. As $X$ is isomorphic to $\mathcal{S}^{n}$, it correspondingly has a subgroup $Q_{X}$ isomorphic to $Q^{n}$, and containing $Z_{X}$. We thus have a chain of subgroups $Z_{X}\leq Q_{X}\leq X\leq G$. Since $Z_{X}$ detects SWCs of $G$, we infer that $Q_{X}$ also detects the SWCs of $G$.
@@ 3.4. Gow’s Formula
Theorem 3.8 ([gow1985real], Theorem 1).
Let $\pi$ be an irreducible self-dual representation of $G$ with central character $\omega_{\pi}$ and Frobenius-Schur Indicator $\varepsilon(\pi)$. Then, we have
@@
We simply call this equality Gow’s formula. For irreducible orthogonal $\pi$ of $G$, it means $\chi_{\pi}(\mathbbm{1})=\chi_{\pi}(-\mathbbm{1})$. This is same as $\chi_{\pi}(g_{0})=\chi_{\pi}(g_{n})$, which leads to
@@
[…]
@@
[…]
In the formula above, $\left\lceil\cdot\right\rceil$ denotes the ceiling function; in particular
$\left\lceil\frac{n-2}{2}\right\rceil=\begin{cases}\frac{n-2}{2},&\text{if $n$ is even},\\[6.0pt] \frac{n-1}{2},&\text{if $n$ is odd}.\end{cases}$
@@ 4.1. The case $n=1$
When $\pi$ is irreducible orthogonal, $m_{1}(\pi)=0$ by Corollary 3.8.1. Therefore, $w(\pi)$ is trivial for such representations. (Compare [NSSL2, Theorem 1.2].)
@@ 4.2. The case $n=2$
[…]
@@
Next is a representation induced from $X$. Let $\psi$ be an odd linear character of $\mathbb{F}_{q^{2}}^{\times}$ with $\psi^{q}\neq\psi$. According to [fulton, Section 5.2], there is an irreducible cuspidal representation $\sigma_{\psi}$ of $\mathcal{S}$ corresponding to $\psi$, with degree q1q-1. Viewing $X$ as the product $\mathcal{S}\times\mathcal{S}$ as in Section 3.1, we form the external tensor product representation $\sigma_{\psi}\boxtimes 1$ of $X$. Finally, let $\pi_{2}=\operatorname{Ind}_{X}^{G}(\sigma_{\psi}\boxtimes 1)$. Here, $\deg(\pi_{2})=q^{2}(q-1)(q^{2}+1)$ and again as $\psi$ is odd, $\chi_{\pi_{2}}(-\mathbbm{1})=-\deg\pi_{2}.$ Also, $\chi_{\pi_{2}}(g_{1})=0$: observe that the only conjugates of $g_{1}$ in $X$ are $\pm g_{1}$ and there is a Weyl element sending $g_{1}$ to $g_{1}$ upon conjugation. (And again use the Frobenius character formula.)
@@ 4.4. Weil Representations
defines a group isomorphism $\mathbb{F}_{q}\overset{\sim}{\to}\widehat{\mathbb{F}}_{q}$. Let $\alpha$ be a non-square, and $1$ be the identity in $\mathbb{F}_{q}^{\times}$. According to Gérardin [PGWeil, Theorem 2.4(d)], the symplectic group $G=\operatorname{Sp}(2n,q)$ has two nonisomorphic Weil representations $\mathcal{W}_{\lambda_{1}}$, $\mathcal{W}_{\lambda_{\alpha}}$ associated to the characters $\lambda_{1}$, $\lambda_{\alpha}$. Set $\mathcal{W}=\mathcal{W}_{\lambda_{1}}$ and $\mathcal{W}^{\prime}=\mathcal{W}_{\lambda_{\alpha}}$. Both $\mathcal{W}$, $\mathcal{W}^{\prime}$ are complex representations of degree $q^{n}$, and decompose into two irreducible representations as
@@
From[PGWeil, Corollary 4.8.1], the character values of $\mathcal{W}$ at $g_{i}$ are given by:
@@
where $f_{k}]_{i}$ is the coefficient of $x^{i}$ in $f_{k}(x)=(1-x)^{k}(1+x)^{n-k}$. Note that $f_{k}]_{n-i}=(-1)^{k}[f_{k}]_{i}$ for each $i=0,\ldots,n$, so that
@@
We have already shown this for $\mathcal{W}$. In [PGWeil], we see that $\chi_{\mathcal{W}}(g_{i})=\chi_{\mathcal{W}^{\prime}}(g_{i})$ for each $i$. Thus, the formula also holds for $w(S(\mathcal{W}^{\prime}))$. ∎
@@ 5. Universal SWCs
This is implicit in [FP, Theorem 6.1, page 283]. ∎
@@
By Lemma 3.2, it is enough to show that the restriction
$\operatorname{H}^{i}(X_{m})^{S_{m}}\to\operatorname{H}^{i}(X_{n-1})^{S_{n-1}}$ is injective. Write $R_{n}=P_{n}\otimes_{\mathbb{F}_{2}}Q_{n}$, where $P_{n}=\mathbb{F}_{2}[\mathcal{E}_{1},\ldots,\mathcal{E}_{n}]$ is the subalgebra generated by $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$ and $Q_{n}=\mathbb{F}_{2}[\mathcal{F}_{1},\ldots,\mathcal{F}_{n}]$ is the subalgebra generated by $\mathcal{F}_{1},\ldots,\mathcal{F}_{n}$ in $\operatorname{H}^{*}(\mathcal{S}^{n})$. From above, we need to show that the restriction of the map $P_{m}\otimes Q_{m}\to P_{n-1}\otimes Q_{n-1}$ to the degree $k$ part is injective for all k<4n1k<4n-1.
@@
If $i^{*}_{n}(\alpha)=0$, then all $c_{\mathbf{r},\mathbf{s}}$ are zero since $\mathcal{E}^{\mathbf{r}}\mathcal{F}^{\mathbf{s}}$ are linearly independent. This implies $\alpha=0$, and hence we have the injectivity of $\iota^{*}_{n}$ for k<4n1k<4n-1. ∎
@@ A.1. The Polynomials $\operatorname{P}_{m,n}$
It is symmetric in the $x_{i}$ and the $y_{j}$ separately, so by the fundamental theorem of symmetric polynomials [b47, Chapter 4, Section 6, Theorem 1], there is a unique polynomial $\operatorname{P}_{m,n}\in\Lambda$ such that
@@
A.2. Characteristic Classes of External Tensor Products
@@
We will sketch the proof for SWCs in this Appendix; the case of CCs is similar. Iterating Proposition A.1 gives:
Corollary A.1.1.
@@
A.3. Products of Vector Bundles
@@
Proposition A.2 ([milnor], Chapter 7, Problem 7-C).
@@
Consider the projection maps $p_{i}:B_{1}\times B_{2}\to B_{i}$. Let $p_{i}^{*}E_{i}$ be the pullback of $E_{i}$ by $p_{i}$ consisting of elements $((b_{1},b_{2}),e_{i})\in(B_{1}\times B_{2})\times E_{i}$ such that $\Pi_{i}(e_{i})=p_{i}(b_{1},b_{2})=b_{i}$. These are vector bundles over the same base space $B_{1}\times B_{2}$. Thus, we construct their internal tensor product, and put $E_{1}\boxtimes E_{2}:=p_{1}^{*}E_{1}\otimes p_{2}^{*}E_{2}$, which is again a vector bundle over $B_{1}\times B_{2}.$ This with the naturality of SWCs and Proposition A.2 gives:
Proposition A.3.
@@
A.4. Vector Bundles From Representations
Let $G$ be a finite group, and $(\rho,U)$ be a real representation of $G$. Associated to $G$ is a classifying space $BG$ with a contractible right principal $G$-bundle EGEG. From $(\rho,U)$ one can form the associated real vector bundle $EG[U]$ over $BG$. To define $EG[U]$, first form the product $EG\times U$; it is $G$-space under the action $g\cdot(x,u)=(x\cdot g^{-1},\rho(g)u)$. Then $EG[U]$ is the quotient of $EG\times U$ by this action. Then put $w_{i}^{\mathbb{R}}(\rho)=w_{i}(EG[U])$. (See for instance [Benson] or [GKT].) The singular cohomology $\operatorname{H}^{*}(BG,A)$ is naturally isomorphic to the group cohomology $\operatorname{H}^{*}(G,A)$ for any abelian group $A$.
Let $(\pi,V)$ be a complex orthogonal representation of $G$. There is a representation $(\rho,U)$, with $U$ a real vector space, so that $\rho\otimes_{\mathbb{R}}\mathbb{C}\cong\pi$ (see [BrokerDieck, Chapter II, Section 6] for instance). Such a representation is called a real form of $\pi$; it is unique up to isomorphism. We now can define $w(\pi):=w^{\mathbb{R}}(\rho)$.
@@ A.4. Vector Bundles From Representations
[…]
where $\mathcal{U}_{i}=EG_{i}[U_{i}]$. We may identify $BG_{1}\times BG_{2}$ with $BG$. Now, Proposition A.1 for SWCs follows from Proposition A.3 with $E_{i}=\mathcal{U}_{i}$.
Acknowledgments.
Part of this paper comes out of the first author’s Ph.D. thesis [Malik.thesis] at IISER Pune, during which she was supported by a Ph.D. fellowship from the Council of Scientific and Industrial Research, India. Both authors thank the Chennai Mathematical Institute for its hospitality during many visits.